Suppose the market demand curve for a good is Qd=1000-10p and the market supply curve is given by Qs=100+20p required.

 

i)  Calculate the equilibrium price. (4 Marks)

ii) Calculate the equilibrium quantity. (2 Marks)

iii) Illustrate the concept of market equilibrium. (4 Marks)


1
Expert's answer
2021-10-27T13:55:34-0400

Solution:

i.). At equilibrium:

Qd = Qs

1000 – 10p = 100 + 20p

1000 – 100 = 20p + 10p

900 = 30p

P = 30

Equilibrium price = 30

 

ii.). Substitute in the demand function to derive the equilibrium quantity:

Qd = 1000 – 10p

Qd = 1000 – 10(30)

Qd = 1000 – 300 = 700

Qd = 700

Equilibrium quantity = 700

 

iii.). A market is in equilibrium when the quantity demanded at the market price equals the quantity supplied. The equilibrium price or market clearing price is the price at which the quantity demanded equals the quantity supplied, and the corresponding quantity is the equilibrium quantity.

 

This is depicted by the below graph: